Previous | Calendar |
October |
2012 |
Next | ||
---|---|---|---|---|---|---|
Sunday | Monday | Tuesday | Wednesday | Thursday | Friday | Saturday |
30JD274 |
1JD275 |
2JD276 |
3JD277 |
4JD278 |
5JD279First snowfall of the season (1"?). |
6JD280lost network to tower (short power outage in afternoon) |
7JD281 |
8JD282at approx 11 MST, network restored Columbus Day |
9JD283 |
10JD284 |
11JD285 |
12JD286Rain (in late afternoon). |
13JD287 |
14JD288 |
15JD289 |
16JD290Site Visit (SpB): Details , Photos (21) Span cylinder at 350 psi. N2 at 850 psi from 7:57 - 13:17 MST, russter2 removed and hard drive replaced at 15:10 MST, swapped battery for soil logger (voltage too high) at 15:32 MST, adjusted flow in 6262 cal gases Windy and cool (feels like winter is coming). |
17JD291Strong winds. Snowed during night of 10/16. |
18JD292lost contact with phenoCam (or possible camera hardware problem?) |
19JD293 |
20JD294Site Visit (HA, GS): Details Measure soil respiration and collect sap flow data at N. Canopy Tower (HA, GS) at approx 11 MST, ATI K-probe stops working |
21JD295 |
22JD296 |
23JD297Site Visit (SpB): Details , Photos (32) Span cylinder at 300 psi. N2 at 690 psi replaced UPS battery at base of tower cleaned up / painted the MRC snow temp probes at 10:30 MST, swapped battery charger for soil logger (voltages still too high) at 12:10 MST, re-connected grounding wire for ATI K-probe (starts working again) at 13:45 MST, changed orientation of 2.5m LI-7500, re-leveled 2.5m REBS Warm and sunny (clouded up in afternoon). |
24JD298Significant snowfall (5-6"?). |
25JD299 |
26JD300 |
27JD301 |
28JD302 |
29JD303at 11am, meet with JK,SM,PB (Full "Hunters" Moon) |
30JD304 |
31JD305Site Visit (SpB): Details , Photos (36) Span cylinder at 250 psi. N2 at 400 psi from 11:35-12:45 MST, maintenance on the LI-7200 (cleaned cell, replaced dessicant/scrubber bottles) at 12:50 MST, swiped krypton hygrometer at 15:50 MST, started ops60 and ucb_16.csi; added two Tcs to UCB logger. Warm, sunny, calm (snow on ground). Halloween! |
1JD306 |
2JD307 |
3JD308at 00:50 MST, ATI stops working site visit by Noah Molotch bioforestmet class |